OpenStudy (anonymous):

 Simplify (a^3b c^3)^2 * (b^5c)^3

sam (.sam.):

(a^(3)bc^(3))^(2)*(b^(5)c)^(3) a^(6)b^(2)c^(3*2)*(b^(5)c)^(3) a^(6)b^(2)c^(6)*(b^(5)c)^(3) a^(6)b^(2)c^(6)*b^(15)c^(3) Multiply a^(6)b^(2)c^(6) by b^(15)c^(3) to get a^(6)b^(17)c^(9). a^(6)b^(17)c^(9)

OpenStudy (anonymous):

\[(a^3b c^3)^2 \times (b^5c)^3\]\[(a^6b c^6) \times (b^{15}c3)\]\[(a^6b^{1+15} c^{6+3})\]\[(a^6b^{16} c^{9}) = Answer\]